20. a) The number of accidents occurring in a city in a day is a Poisson variate with mean 0.8. Find the probability that on a randomly selected day (i)There are no accidents (ii)There are accidents
b) The incidence of an occupational disease in an industry is such that the worker have 25% change of suffering from it. What is the probability that out of 5 workers, at the most two contract that disease?
1
Expert's answer
2019-10-23T14:56:27-0400

a)

X = the number of accidents occurring in a city in a day.


"P(X=x)={\lambda^x e^-\lambda \over x!}, x=0,1,2,..."

(i) Find the probability that on a randomly selected day there are no accidents

"1\ day=>\lambda=0.8"


"P(X=0)={0.8^0 e^{-0.8} \over 0!}=e^{-0.8}\approx0.449329"

(ii) Find the probability that on a randomly selected day there are accidents


"P(X>0)=1-P(X=0)=1-{0.8^0 e^{-0.8} \over 0!}=1-e^{-0.8}\approx""\approx0.550671"

b)

X = the number of diseased workers.

Binomial distribution


"b(x;n, p)=\binom{n}{x}p^x(1-p)^{n-x},x=0,1,2,...,n"

Given that "n=5,p=0.25."

Find the probability that out of 5 workers, at the most two contract that disease


"P(X\leq2)=P(X=0)+P(X=1)+P(X=2)="

"=\binom{5}{0}0.25^0(1-0.25)^{5-0}+\binom{5}{1}0.25^1(1-0.25)^{5-1}+"

"+\binom{5}{2}0.25^2(1-0.25)^{5-2}="

"=0.75^5+5(0.25)(0.75)^4+10(0.25)^2(0.75)^3="

"=0.2373046875+0.3955078125+0.263671875="

"=0.896974375"


The probability that out of 5 workers, at the most two contract that disease is "0.896974375"